OpenStudy (anonymous):

A wooden pyramid, 6 inches tall, has a square base. A carpenter increases the dimensions of the wooden pyramid by a factor of 3 and makes a larger pyramid with the new dimensions. Describe in complete sentences the ratio of the volumes of the two pyramids. i know the ratio is 1/6 but im not sure how to explain everything??

OpenStudy (anonymous):

?? can anyone help me

OpenStudy (anonymous):

the ratio of the linear measurements is 1:3 the ratio of volumes can be found by raising each value in the ratio to the power of 3 (1)^3 :(3)^3 = 1:27 a test.... height 6 and assume the square base has side length of 1 inch V=1/3×12×6 V = 2 inches^3 now with the dimensions tripled h = 18 and square side length is 3 V=1/3×32×18 V = 54 inches ^3 the ratio of the volumes is 2:54 or 1: 27

OpenStudy (anonymous):

thnk u

OpenStudy (anonymous):

no i i get it

OpenStudy (anonymous):

how do u get 1/3

OpenStudy (anonymous):

It's increasing by a factor of three... meaning that it's growing three times bigger not six. Therefore, for every one dimension it's increasing by itself 3 times. Make sense?

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

so like multiplying right

OpenStudy (anonymous):

kind of...

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

how do you know whaT to multiply by when ur calculating everything

OpenStudy (anonymous):

Well I'm going to answer your question with a question.. How did you get 6?

OpenStudy (anonymous):

i had guessed because you were starting of with 6 inches

OpenStudy (anonymous):

The way I think about it is basically if you started at 1 inch and it grew 3 more inches then it would be a ratio of 1/3 Since it starts with 6 and grow three times it would be 6/18 and then you can simplify that into 1/3

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

but i thought that u said it was 1/3

OpenStudy (anonymous):

so 1/3 is just the simplified ratio?

OpenStudy (anonymous):

no. if someone were to look at this and have to figure out the ratio without any math they would easily get 1/3. I see where you got 1/6 though.

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

Get it now?

OpenStudy (anonymous):

yes lol i was so confused

OpenStudy (anonymous):

thank u so much

OpenStudy (anonymous):

Anytime! :)

OpenStudy (anonymous):

wait when you were mutpliying 1/3x12x6 how did u get the 12